OpenStudy (anonymous):

@wio pretty tiered so im not sure why this isn't adding up in my head but.. y=2x^2-4x+1 .. i can only see as (2x-1)(x-1).. and that gives 2x^2-3x+1..

OpenStudy (psymon):

Yeah, youre not going to get an answer that clean.

OpenStudy (psymon):

Need quadratic formula for it, if you know the formula.

OpenStudy (anonymous):

\[ -2x-x\neq -4x \]

OpenStudy (anonymous):

oh the -b= the square root of (x1-x2)^2-(y1-y2)^2 ?

OpenStudy (psymon):

Not quite. \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Can't you use quadratic formula to find the roots here?

OpenStudy (anonymous):

hmm, i don't remember discussing that formula this week. I know it. so i guess ill use it.

OpenStudy (anonymous):

i got rad 2 over 2..

OpenStudy (psymon):

So....(4 +/- sqrt(16-8))4 \[\frac{ 4\pm2\sqrt{2} }{ 4 }=1\pm \frac{ \sqrt{2} }{ 2 } \] Somehow you got rid of your +/- and the 1. You need the plus and minus for your solutions, not just the one sqrt2/2 solution.

OpenStudy (anonymous):

oh yea. Srry. Just thought the 1 wasn't important. Im starting to remember this stuff.

OpenStudy (psymon):

It's all important. You want the correct answer, not something kind of close xD

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (psymon):

Yep, sure :3